# Layer-Wise Quantization: A Pragmatic and Effective Method for Quantizing LLMs Beyond Integer Bit-Levels

### v2

@@ 3.2.2 GPT-Q
GPT-Q is a post-training quantization technique specifically designed for GPT models Frantar et al. (2023). Utilizing a dataset, it calculates the necessary scaling factors for quantization. After training, GPT-Q assesses the distribution of weights using this dataset to determine optimal scaling factors for converting floating-point representations to lower-bit formats such as int8 or int4. We only have several data points for GPT-Q and it adds significant execution time overhead to our experiments. Both Quanto and GPT-Q are available in HuggingFace library Wolf et al. (2019). We modified it for our use-case in the same way as described for Quanto.
@@ 4.1 Models and Hyperparameters
We studied quantization on 5 LLMs from different model families and different sizes – LLaMa-2-7b Touvron et al. (2023), LLaMa-2-13B, Mistral-7b Jiang et al. (2023), QWEN-1.8b, and QWEN-7b Bai et al. (2023). We evaluated these LLMs on 8 A100 GPUs with a batch size of 1 for inference using LLM harness library Gao et al. (2023).
@@ 4.2 Evaluation Datasets
We select five diverse NLP tasks for evaluating the quantization effects: Winogrande Sakaguchi et al. (2021), ARC-easy Clark et al. (2018), PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), and MMLU Hendrycks et al. (2020). We also evaluate our approach on two generation datasets to cover diverse tasks of reasoning and answer generation: GSM8K Cobbe et al. (2021), which contains math questions, and the Natural Questions (open) dataset Kwiatkowski et al. (2019), which consists of open-domain answer generation task.
@@ 5 Discussion of Results
Our main results are shown in Figure 2 and 3 (please also see Figure 8, 9, and 10 in Appendix). In both these figures, we ranked the layers in the respective LLMs in descending order of their importance score. The left most point in the plots indicates that all layers of Mistral-7b and LLaMa2-7B are quantized in 4-bits; as we move to the right on x-axis, we quantize the next least important layer to 2 bits. For example, the overall bit size of 3.75 mentioned on the x-axis represents 28 (most important) layers in 4-bit and 4 (least important) layers in 2-bit. The horizontal red dotted lines indicate the performance of the entire model represented in 8bit precision; as shown, this performs very similarly to the full model in 4 bit precision (the top left most point in Figure 2 and 3). Because the gap between the full model in 8 bit vs. 4 bit precision was less than 1% across the majority of the datasets, we focus mostly on quantizing below 4-bit precision in our experiments.
4 to 2 bit 8 to 4 bit Models Layers GSM8K NQ_open GSM8K NQ_open 2 bits F1 F1 LIM Ordering LLama-Ins 5 7.5 15.0 10.5 36.6 10 1.5 6.7 13.5 37.9 15 0.5 3.1 11.0 35.4 Mistral-Ins 5 24.5 16.2 34.5 29.6 10 20.0 9.2 36.5 29.8 15 4.5 5.1 34.0 27.7 Z Ordering LLama-Ins 5 6.0 11.7 12.5 37.3 10 3.5 5.3 12.5 36.9 15 1.0 1.8 10.0 34.7 Mistral-Ins 5 25.0 15.4 37.0 28.5 10 10.5 8.7 34.5 30.1 15 1.0 2.2 34.0 29.3
@@
    Variable quantization is useful: The first key finding of our work is that a fixed quantization technique can be extended to a variable number of bits by quantizing different layers at different bits according to their importance. This allows LLMs to be fit in the exact memory requirement of the user while retaining more of the original performance. Overall, our method of layer-wise quantization, guided by layer importance, proves to be an efficient strategy for attaining adaptable precision bit levels.
@@
    Layer importance scoring is crucial: In the figures we compare layer ranking using our LIM and ZD importance scores with ranking using a reverse LIM and random ordering. As seen in fig. 2 and 3 the quantization of least important layers from 4-bit to 2-bit as per the LIM score ranking shows strong performance retention. In contrast, quantizing based on the reverse of LIM score shows much worse performance when most important layers are quantized to 2-bit, highlighting the strength of meaningful layer ranking. Further, LIM and ZD ranking performs substantially better than random ordering of layers baseline where we quantize layers to lower bits randomly. Lastly, LIM performs better than ZD, but the differences are not large. We consider this a success for ZD, which is simpler and does not require calibration data.
@@
    Layer-wise Quantization is useful until 3.0-3.25 bits: As shown in fig. 2 and 3, our first key observation is that quantization to 8 bit barely affects performance (red vs. purple line in each plot). While there is marginal drop in performance when models are quantized to 4 bits, we observed really noticeable drops only after the models are dropped below 3.0-3.25 bits on average using Quanto. For example, the bit size for which performance drops below 90% on Winogrande for Mistral-7b, LLaMa-7b, QWEN-7b, and LLaMa2-13b are 3.2, 3.1, 3.85, and 2.85 respectively.
@@
    Figure 5: Comparison of LLaMa2-7b quantized between 8 and 4 bits with LLaMa2-13b quantized between 4 and 2 bits to check when the performance intersects
@@
    Our method is applicable to different quantization techniques: Tables 1 and 2 summarize the overall results when quantizing individual layers with Quanto and GPT-Q, respectively. The tables show that our method can be coupled with any other quantization techniques. On average, GPT-Q leads to an average of 4% better accuracy than Quanto across all 5 tasks. Additionally, GPT-Q enables models such as LLaMa2-13B to be quantized down from 4 bits to 3.25 bits with less than a 3% loss in average accuracy, as seen in Table 2. Importantly, similar to Quanto and GPT-Q, our methodology of layer-wise quantization can be extended to any quantization technique.
@@
    Effect on generation tasks: — We also evaluate our approach of variable quantization of different layers on generation tasks. As shown in table 3, we observe substantial drop in performance on both GSM8K and NQ_open generation tasks when quantizing more layers in 2 bits. Importantly, the performance drop in these generation tasks is more drastic when compared to the average performance drop in classification tasks (table 1), emphasizing the need for more dedicated research in quantization for generation tasks.
@@ 6 Analyses
We compare our variable quantization against variable pruning (using the same layer importance ranking) as an alternative for the same goal of reducing model memory requirement. In Figure 4(a), we show that for not extreme quantization levels, it is significantly better to move layers into lower quantization levels instead of removing them. For example, when 2 least important layers are removed resulting in remaining 30 layers (each in 8 bit) of LLaMa2-7b, the average performance drops to 62.7% (shown by red line denoting work by Gromov et al. (2024)). But on the quantization counterpart with same memory i.e., when 4 layers are quantized to 4 bit and remaining 28 layers are in 8 bit (shown by blue line), performance remains intact close to 66.8% as shown in fig. 4(a). When 12 layers are removed, the performance drops around 53% on average while having 8 layers in 8 bits and 24 layers quantized to 4 bits (shown by blue curve) to maintain the same size, average performance still remains intact and close to 66%. This highlights the important finding that quantization until 4 bits overall is a substantially more effective strategy compared to pruning for model compression.
On the other hand, in case of extreme levels of quantization (i.e., $<4-bits$) as shown fig. 4(b), it is better to plainly remove layers. For example, as shown in fig. 4(b) where the model is initially quantized with 4-bits, removing $>7$ layers results in better average performance when compared to quantizing $>14$ layers in 2 bits. Thus, when model compression is requires to be the equivalent of $<3bits$, pruning maybe the more effective strategy.
@@
We further evaluate and compare feasibility of quantizing larger LLMs more drastically (i.e., $<4bits$ by quantizing less important layers in 2 bits and keeping more important in 4 bits) or quantizing smaller LLMs moderately (i.e., $<8bits$ by quantizing less important layers in 4 bits and keeping more important in 8 bits). As shown in fig. 5, we quantize LLaMa2-13B $<4bits$ and LLaMa2-7B $<8bits$ and compare them across different memory sizes. Our emperical finding suggests its more beneficial to quantize larger LLMs to smaller bits but only until a certain point, after which layer-wise quantizing smaller LLMs moderately ($6~{}bits$) shows better performance.
@@ 7 Conclusion
We introduced a simple, flexible quantization approach that quantizes different layers at different bits based on their importance. We presented two layer importance scoring techniques which when used to select more important layers for quantizing them in 4 bits and less important layers in 2 bits lead to strong performance retention across several LLMs even until 2.85 overall bit size. Our work presents several key practical findings such as layer-wise quantization is more effective for larger LLMs (with more number of layers), in the same memory setting; quantization is better than pruning until a certain bit precision level, text generation tasks are affected more with quantization, etc. Overall, our work introduces layer-wise quantization and presents detailed empirical findings for motivating future research in this direction.
8 Limitations
Limitations of our work are as follows:
  • Our experiments and results have focused more on quantization to lower bits i.e., $<4bits$ to highlight gains from layer-wise quantization based on their importance. As mentioned in section 5, we observed very similar performance between fp16 and 8-bit quantized model, minimal drop ($<$ 1% on average) in performance between 8-bit and 4-bit quantized model (please see table 4 in appendix). We did not observe meaningful changes in performance from layer-wise quantization between 8-bit and 4-bit because of such minute difference between their performances. Our study is limited to two level quantization i.e., more important layers in 4 bits and less important layers in 2 bits. One can also potentially apply three level of quantization i.e., most important layers in 8 bits, moderately important in 4 bits and least important in 2 bits. In our experiments, we observed three level quantization to be always worse than two level quantization for similar memory sizes (please see section A.1 in Appendix). .
  • We have proposed LIM and ZD scores for rating importance of each layer in LLM. These scores are very simple to implement, and only LIM needs a calibration corpus which is also an (readily available) unlabelled corpus. As future works, there can be more sophisticated measures to find importance of layers. For example, supervised methods that can show layer-wise impact on specific annotated NLP tasks Zhu et al. (2023a), in depth model interpretability approaches Singh et al. (2024); Sun et al. (2023), analyses of block of layers and their interactionsYang et al. (2024), etc. In this work, we limited and focused our study on LIM and ZD score to emphasize more on the main contribution of the paper - quantizing different layers at different bits as per layer importance. To the best of our knowledge, our work is the first to introduce variable quantization layer wise as per their importance and our study can be easily extended with more layer importance measuring scores in future work.
  • We presented comparison between pruning and quantization in section 6.1. A potential setup can also be combining pruning and quantization by removing few of the least important layers, quantizing remaining less important layers in lower 2 bits and keeping more important layers in 8 bits or above. Our work is primarily focused on introducing the idea of achieving variable quantization by quantizing different layers as per their importance. We leave exploration of combining quantization and pruning for future works.
9 Ethical Considerations
Our study is focused on post training quantization (PTQ) of LLMs. We have not finetuned LLMs for any specific task or data and have selected well established LLMs such as LLaMa, Mistral, and QWEN in our experiments. Our study also uses widely used evaluation datsets such as MMLU, HellaSwag, ARC easy, etc. that contain safe test cases. Thus, we believe our experiments do not contain any harmful cases and to the best of our knowledge, we did not observe any unsafe outputs in our evaluations.
@@ Appendix A Appendix
Figure 6: We first rank the layers in an LLM (e.g., LLaMa2-13B) using an importance score (shown here is ranking based on our Layer Input Modification (LIM) score, see section 3.1).
@@ A.1 Quantizing Layers Using 3 Levels
In all of our experiments, we have focused on two level quantization, i.e., either less important layers are quantized in 2 bits and more important layers in 4 bits or less important in 4 bits and more important in 8 bits. As a plausible variant, LLM layers can also be easily quantized using three levels, i.e., least important layers in 2 bits, moderately important in 4 bits, and the most important ones in 8 bits. In our study, we observed three level quantization almost always performs worse than two level quantization. We show three level quantization of LLaMa2-7b with fixed overall model bit size of 4bit. We first quantize all the 32 layers in 4 bits as shown by the leftmost bar of section A.1. We then convert two least important layers to 2 bits each and one most important layer to 8 bit, thus maintaining the overall bit size of the model to 4 bit. This is represented by second bar from the left in fig. 7. We repeat the same process of converting two more layers in 2bits and a more important layer to 8bit represented by the consecutive bars in fig. 7. As observed, three level quantization always performs worse than one level quantization when the target bit-level is the same, thus we don’t propose the technique as a way to achieve better performance for a set quantization level, but as a way to achieve a variable level of quantization while retaining maximum performance.
@@
Figure 8: Similarly to the other bit plots the graphs showcase the accuracy on four distinct data sets when quantizing LLaMa2-7b from full 4-bit quantization to 2-bit by moving less important layers in 2-bit quantization
@@
Table 4: Accuracy results of different models across various tasks for 8bit and 4bit quantization.

### v3

@@ 3.3.2 GPT-Q
GPT-Q is a post-training quantization technique specifically designed for GPT models (Frantar et al., 2023). Utilizing a dataset, it calculates the necessary scaling factors for quantization. After training, GPT-Q assesses the distribution of weights using this dataset to determine optimal scaling factors for converting floating-point representations to lower-bit formats such as int8 or int4. We only used a few data points for GPT-Q as it adds significant execution time overhead to our experiments. We modified the GPT-Q implementation in the Hugging Face library (Wolf et al., 2019) for our use-case in the same way as described above for Quanto.
@@ 4.1 Models and Hyperparameters
We studied quantization on 5 LLMs from different model families and different sizes – LLaMa-2-7b (Touvron et al., 2023), LLaMa-2-13B, Mistral-7b (Jiang et al., 2023), QWEN-1.8b, and QWEN-7b (Bai et al., 2023). We evaluated these LLMs on 8 A100 GPUs with a batch size of 1 for inference using an LLM harness library (Gao et al., 2023).
@@ 4.2 Evaluation Datasets
We select five diverse NLP tasks for evaluating the quantization effects: Winogrande (Sakaguchi et al., 2021), ARC-easy (Clark et al., 2018), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), and MMLU (Hendrycks et al., 2020). We also evaluate our approach on two generation datasets to cover diverse tasks of reasoning and answer generation: GSM8K (Cobbe et al., 2021), which contains math questions, and the Natural Questions (open) dataset (Kwiatkowski et al., 2019), which consists of open-domain answer generation task. To calibrate the LIM score we used 50 samples from PG19 (Rae et al., 2019b), a data-set that is comprised of books published before 1919.
@@ 5 Discussion of Results
Our main results are shown in Figure 2 and 3 (please also see Figure 8, 9, and 10 in Appendix). In both these figures, we ranked the layers in the respective LLMs in descending order of their importance score. The left most point in the plots indicates that all layers of Mistral-7b and LLaMa2-7B are quantized in 4-bits; as we move to the right on x-axis, we quantize the next least important layer to 2-bits. For example, the overall bit size of 3.75 mentioned on the x-axis represents 28 (most important) layers in 4-bit and 4 (least important) layers in 2-bit. The horizontal red dotted lines indicate the performance of the entire model represented in 8-bit precision; as shown, this performs very similarly to the full model in 4-bit precision (the top left most point in Figure 2 and 3). Because the gap between the full model in 8-bit vs. 4-bit precision was less than 1% across the majority of the datasets, we focus mostly on quantizing below 4-bits precision in our experiments.
4 to 2 bits 8 to 4 bits Models Layers GSM8K NQ_open GSM8K NQ_open 2-bits F1 F1 LIM Ordering LLama-Ins 5 7.5 15.0 10.5 36.6 10 1.5 6.7 13.5 37.9 15 0.5 3.1 11.0 35.4 Mistral-Ins 5 24.5 16.2 34.5 29.6 10 20.0 9.2 36.5 29.8 15 4.5 5.1 34.0 27.7 Z Ordering LLama-Ins 5 6.0 11.7 12.5 37.3 10 3.5 5.3 12.5 36.9 15 1.0 1.8 10.0 34.7 Mistral-Ins 5 25.0 15.4 37.0 28.5 10 10.5 8.7 34.5 30.1 15 1.0 2.2 34.0 29.3
@@
Method Layers / Threshold Avg. Bits Wikitext Perplexity C4 Perplexity
       Our Method 32 3.94 5.6134 7.5941
27 3.90 5.6439 7.6101
22 3.85 5.6597 7.6326
17 3.81 5.6631 7.6487
12 3.76 5.6848 7.6737
6 3.71 5.7009 7.6938
0 3.65 5.7145 7.7126
       SpQR 1.00% 3.94 5.6134 7.5941
0.85% 3.90 5.6403 7.6523
0.70% 3.85 5.6929 7.6878
0.55% 3.80 5.6883 7.7041
0.40% 3.75 5.7054 7.7060
0.25% 3.71 5.7151 7.7140
0.10% 3.65 5.7145 7.7126
Table 4: Comparison of our method and SpQR (Dettmers et al., 2023) on LLaMa2-7B on the same avg. bit level/memory requirement. Our technique aims to optimize performance given memory constraints, allowing quantization to fit more precisely within available memory, which, as shown, improves results over fixed bit-width quantization. For SpQR, we fixed the bit levels to 3 and beta values to 4, varying the outlier threshold from 0.1% to 1%. For our method, we applied SpQR with thresholds of 0.1% to less important layers and 1% to more important layers based on LIM ordering. The second column indicates the number of layers with a higher threshold for our method and the outlier threshold for SpQR. We bold the higher values and we use italics to denote identical values. Note that each row in the “Our Method” block should be compared with the row at the same position in the “SpQR” block. That is, our setting with 32 layers in higher precision is comparable to SpQR with the 1.00% outlier threshold; our setting with 0 layers in higher precision is comparable with SpQR with all layers using 0.10% outlier threshold.
@@
    Variable quantization is useful: The first key finding of our work is that a fixed quantization technique can be extended to a variable number of bits by quantizing different layers at different bits according to their importance. This allows LLMs to retain more of the original performance while fitting in a reduced memory budget. Overall, our method of layer-wise quantization, guided by layer importance, proves to be an efficient strategy for attaining adaptable precision bit levels.
@@
    Layer importance scoring is crucial: In the figures we compare layer ranking using our LIM and ZD importance scores with ranking using a reverse LIM and random ordering. As seen in fig. 2 and 3 the quantization of least important layers from 4-bit to 2-bit as per the LIM score ranking shows strong performance retention. In contrast, quantizing based on the reverse of LIM score shows much worse performance when most important layers are quantized to 2-bit, highlighting the strength of meaningful layer ranking. Further, LIM and ZD ranking performs substantially better than random ordering of layers baseline where we quantize layers to lower bits randomly. Lastly, LIM performs better than ZD on average, but the differences are not large. In some cases, i.e., Figure 9, ZD performs considerably better. We conjecture this is due to the fact that for this LLM the information gathered from the dataset used to calibrate LIM transferred less well to the evaluation datasets. All in all, we consider this a success for ZD, which is simpler and does not require calibration data.
@@
    Improving dynamic quantization techniques: Table 4 shows that we can use our technique to further improve quantization techniques that are inherently dynamic. Although this is not the main goal of the paper, it further underscores its usefulness and flexibility. The reason why this works is because the SOTA quantization technique SpQR (Dettmers et al., 2023), similarly to most dynamic techniques, choose a fixed percentage of values that will be quantized to higher precision in each layer, while our techniques adds the fact that not all layers are equally important on top of it. This also further shows that our proposed technique can be used an enhancement to any category of quantization techniques.
  4. (4)
    Layer-wise quantization is useful until 3.0–3.25-bits: As shown in fig. 2 and 3, our first key observation is that quantization to 8-bits barely affects performance (red vs. purple line in each plot). While there is marginal drop in performance when models are quantized to 4-bits, we observed really noticeable drops only after the models are dropped below 3.0–3.25 bits on average using Quanto. For example, the bit size for which performance drops below 90% on Winogrande for Mistral-7b, LLaMa-7b, QWEN-7b, and LLaMa2-13b are 3.2, 3.1, 3.85, and 2.85 respectively.
@@
    Our method is applicable to different quantization techniques: Tables 1 and 2 summarize the overall results when quantizing individual layers with Quanto and GPT-Q, respectively. The tables show that our method can be coupled with any other quantization techniques. On average, GPT-Q leads to an average of 4% better accuracy than Quanto across all 5 tasks. Additionally, GPT-Q enables models such as LLaMa2-13B to be quantized down from 4-bits to 3.25-bits with less than a 3% loss in average accuracy, as seen in Table 2.
@@
    Effect on generation tasks: We also evaluate our approach of variable quantization of different layers on generation tasks. As shown in table 3, we observe substantial drop in performance on both GSM8K and NQ_open generation tasks when quantizing more layers in 2-bits. Importantly, the performance drop in these generation tasks is more drastic when compared to the average performance drop in classification tasks (table 1), emphasizing the need for more dedicated research in quantization for generation tasks.
@@ 6 Analyses
Figure 5: Comparison of LLaMa2-7b quantized between 8 and 4 bits with LLaMa2-13b quantized between 4 and 2 bits to check when the performance intersects.
@@
We compare our variable quantization against variable pruning (using the same layer importance ranking) as an alternative for the same goal of reducing model memory requirement. In Figure 4(a), we show that for less extreme quantization levels, it is significantly better to move layers into lower quantization levels instead of removing them. For example, when 2 least important layers are removed resulting in remaining 30 layers (each in 8-bit) of LLaMa2-7b, the average performance drops to 62.7% (shown by red line denoting work by Gromov et al. (2024)). But on the quantization counterpart with same memory i.e., when 4 layers are quantized to 4-bit and remaining 28 layers are in 8-bit (shown by blue line), performance remains intact close to 66.8% as shown in fig. 4(a). When 12 layers are removed, the performance drops around 53% on average while having 8 layers in 8-bits and 24 layers quantized to 4-bits (shown by blue curve) to maintain the same size, average performance still remains intact and close to 66%. This highlights the important finding that quantization until 4-bits overall is a substantially more effective strategy compared to pruning for model compression.
On the other hand, in case of extreme levels of quantization (i.e., $<4-$bits) as shown fig. 4(b), it is better to plainly remove layers. For example, as shown in fig. 4(b) where the model is initially quantized with 4-bits, removing $>7$ layers results in better average performance when compared to quantizing $>14$ layers in 2-bits. Thus, when model compression is requires to be the equivalent of $<3$-bits, pruning maybe the more effective strategy.
@@
We further evaluate and compare feasibility of quantizing larger LLMs more drastically (i.e., $<4$-bits by quantizing less important layers in 2-bits and keeping more important in 4-bits) or quantizing smaller LLMs moderately (i.e., $<8$-bits by quantizing less important layers in 4-bits and keeping more important in 8-bits). As shown in fig. 5, we quantize LLaMa2-13B $<4$-bits and LLaMa2-7B $<8$-bits and compare them across different memory sizes. Our empirical finding suggests it is more beneficial to quantize larger LLMs to smaller bits but only until a certain point, after which layer-wise quantizing smaller LLMs moderately ($\approx 6$-bits) shows better performance.
@@ 7 Conclusion
We introduced a simple, flexible quantization approach that quantizes different layers at different bits based on their importance. We presented two layer importance scoring techniques which when used to select more important layers for quantizing them in 4 bits and less important layers in 2-bits lead to strong performance retention across several LLMs even until 2.85 overall bit size. Our work presents several key practical findings such as layer-wise quantization is more effective for larger LLMs (with more number of layers), in the same memory setting; quantization is better than pruning until a certain bit precision level, text generation tasks are affected more with quantization, etc. Overall, our work introduces layer-wise quantization and presents detailed empirical findings for motivating future research in this direction.
@@ A.1 Quantizing Layers Using 3 Levels
In all of our experiments, we have focused on two level quantization, i.e., either less important layers are quantized in 2-bits and more important layers in 4-bits or less important in 4-bits and more important in 8-bits. As a plausible variant, LLM layers can also be easily quantized using three levels, i.e., least important layers in 2-bits, moderately important in 4 bits, and the most important ones in 8-bits. In our study, we observed three level quantization almost always performs worse than two level quantization. We show three level quantization of LLaMa2-7b with fixed overall model bit size of 4-bits. We first quantize all the 32 layers in 4-bits as shown by the leftmost bar of section A.1. We then convert two least important layers to 2-bits each and one most important layer to 8-bits, thus maintaining the overall bit size of the model to 4-bits. This is represented by second bar from the left in fig. 7. We repeat the same process of converting two more layers in 2-bits and a more important layer to 8-bits represented by the consecutive bars in fig. 7. As observed, three level quantization always performs worse than one level quantization when the target bit-level is the same, thus we don’t propose the technique as a way to achieve better performance for a set quantization level, but as a way to achieve a variable level of quantization while retaining maximum performance.
Figure 6: Visualization of the layer importance score for four different LLMs. Shown here is our Layer Input Modification (LIM) score. The color intensity of each layer, which represents their LIM importance score (darker color indicates higher importance score), highlights that the original layer structure does not have the layers sorted according to their importance.
@@
Figure 8: Similarly to the other bit plots the graphs showcase the accuracy on four distinct data sets when quantizing LLaMa2-7b from full 4-bits quantization to 2-bit by moving less important layers in 2-bits quantization
@@
Table 5: Accuracy results of different models across various tasks for 8bit and 4bit quantization using Quanto as the quantization technique.